OpenStudy (anonymous):

y=1/x+3 What is the vertical asymptote for the above function? x=_____________

OpenStudy (anonymous):

helppp!!! :((

OpenStudy (anonymous):

given, y=1/x+3 y-3=1/x 1/x=y-3 x=1/(y-3)

OpenStudy (anonymous):

-3 is the answer....